# MAPSS: Manifold-based Assessment of Perceptual Source Separation

### v1

@@ 3.4 The Perceptual Match (PM) Measure
Differently from the PS, here the focus shifts to whether an output remains perceptually close to its own reference. Since each cluster of distortions around a reference may be anisotropic, the Mahalanobis distance is again the natural choice. If the output lies well within the bulk of its distortion cluster, the Gamma-tail probability is near one, which may indicate a strong perceptual match. As the output drifts away, the score decays smoothly toward zero, reflecting degradation. Observing Figure 2, in edge cases, when distortions are tightly concentrated, even small mismatches in $a_{i}^{(d)}$ lower PM sharply, while, when distortions are more broadly dispersed, the PM remains tolerant to larger deviations.
The empirical distributions of the frame-level values of the measures are shown in Figure 3. The PM and PS metrics exhibit contrasting distribution patterns. PM values cluster predominantly around zero with minimal density near one, while PS concentrate near one with virtually no occurrence near zero. Although frame-level human speech quality ratings are not publicly available for direct comparison, these patterns raise comparisons to how humans might perceive audio disturbances. The PM distribution aligns intuitively with human perception, as listeners typically penalize speech quality severely when disturbances occur, making ratings near the scale minimum unsurprising. However, real granular human ratings would likely show less extreme clustering around zero due to perceptual and rating scale complexities. The PS behavior presents a more complex interpretative challenge. Previous research suggests that humans perceive leakage as more quality-degrading than self-distortions, particularly in acoustic echo cancellation contexts Khanagha et al. (2024), yet our findings here do not support this hypothesis. Whether this discrepancy stems from dataset characteristics, limitations of the PS measure itself, or the mismatch between granular PS values and aggregated human ratings remains unclear and warrants future investigation beyond the scope of this study.
@@ 4.1 Database
To benchmark the proposed measures, we rely on the Subjective Evaluation of Blind Audio Source Separation (SEBASS) database (Kastner & Herre, 2022), a publicly-available collection of expertly curated listening tests that pools $11,000$ individual human ratings for more than $900$ separated signals, originating from five evaluation campaigns. SEBASS focuses on distinct scenarios of speech mixtures of 4 male and 4 female speakers, separately, with a pair of English and pair of Spanish speakers in each, and music with both a drums and a no-drums subset, each containing 3 music sources. The split between music mixtures with and without drums is crucial, as percussion transients introduce masking patterns that are perceptually and algorithmically distinct from purely harmonic content, thereby stressing leakage and self-distortion behaviors that do not appear in speech mixtures.
Every mixture was fed to 32 different speech separation systems, spanning classic methods such as non-negative matrix factorization (NMF), spatial filtering and modern deep-learning models. Every output utterance of 10 s, sampled at a sampling frequency of 16 kHz, was judged by 10 or more predominantly certified human raters according to the MUSHRA standard Schoeffler et al. (2018). Every rater was asked to provide a grade between 0 and $100$ to the question “grade the basic audio quality of the items under test with respect to the reference signal. Any perceived differences between the reference and the other signal must be interpreted as an impairment”. During grading, the user was exposed to the reference as a baseline and to the hidden reference and mixture anchor as comparative outputs. Figure 4 shows that the speech reference signals have been recorded in a relatively clean environment with signal-to-noise-ratios (SNRs) between 3.9 dB to 41.7 dB, with an average of 25 dB.
The SEBASS dataset suits this study for several reasons. Multilingual coverage of English and Spanish validates language-agnostic behavior, while music tests robustness to highly transient material. Large algorithmic spread creates rich output clusters that stress-test our methodology, and the dense sampling of raters allows for a more reliable estimation of the true mean-opinion score of subjective human opinion. Importantly, SEBASS represents the only available dataset incorporating human quality ratings for source separation tasks. While creating an additional human-rated dataset falls beyond our scope, we would like to note that expanded perceptual data for source separation would significantly benefit and foster future research in this area.
4.2 Pre-processing
We first recognize that English and Spanish speakers rarely participate in the same conversation in real-life scenarios. Thus, we separate each 4-speaker mixture into their English and Spanish speakers, creating for each language two mixtures where the one has a pair of male speakers and the other a pair of female speakers. We acknowledge the bias this contributes, as residuals of English may be present in the output signal of a Spanish speaker, and vice versa. It should be mentioned that listening tests have rendered this cross-language leakage extremely negligible, since as expected, source separation systems are able to leverage languages as a meaningful feature to recognize leakage and remove it.
Every waveform, including references, distortions, and outputs from all sources of the mixture, undergoes independent loudness normalization. We use the EBU Recommendation R-128 European Broadcasting Union (2011) (EBU) and set the target level of each waveform to loudness units relative to full scale (LUFS) of -23. If the peak magnitude of the scaled waveform exceeds one, we attenuate it to avoid digital clipping. This step removes loudness bias, known to wrongly affect both human and algorithmic quality judgments, while preserving inter-speaker level relations across the outputs. Since the PS and PM measures address source separation, we also filter out any frames in which there are not at least two active sources using energy-based thresholding.
When applying diffusion maps, we set $\alpha=1$ in equation (6) to eliminate density-dependent bias from the embedding. This choice ensures that the PS and PM measures reflect the intrinsic geometric structure of the manifold rather than sampling density variations, which would introduce instead artificial distortions into the representation. In the specific case of music with drums, however, we choose to make an exception because density encodes naturally occurring frequency and perceptual salience of transient states, as in drum mixing. Letting $\alpha=0$ retains density bias and makes the PS and PM event-weighted, which we assume aligns with the perceptual experience of the average human listener. We set $t=1$ (11) across the board, to keep the diffusion operator focused on local neighborhoods and not being blurred by multi-step mixing.
Unfortunately, granular subjective human scores are not available to the best of our knowledge, although a protocol to obtain them has been established Hansen & Kollmeier (1999). In fact, for source separation, SEBASS provides these scores at utterance-level resolution of approximately 10 seconds. Since our PM and PS measures operate at finer temporal resolutions of 50 and 10 frames-per-second for speech and music, respectively, then aggregation from frame-level to utterance-level is necessary to compare our measure to human MOS. Based on the characteristic of each measure, we have decided to apply average aggregation for the PM and a more perceptual, PESQ-inspired aggregation for the PS.
We now retrieve the indices used in Section 2 to formalize these aggregations. At trial $l$, let us denote the PM value of the $i$-th output of source-separation system $q$ in time frame $f$ as $\textrm{PM}^{q,l}_{i,f}$. We denote $\mathcal{F}^{l}_{i}\in\mathcal{F}^{l}$ as the subset of time-frame indices in which the $i$-th source is active. Then, the utterance-level PM measure after average aggregation is given by:
Although average aggregation assumes that human listeners perceive global audio quality by weighing local events equally, which is evidently not the case Rix et al. (2001), we chose to carry it for the PM since its behavior already exhibits strong and frequent granular penalties where the score drops to around zero. Thus, it is assumed that standard human behavior that weighs negative experience heavily in the utterance-level score is implicitly carried out by the nature of the PM measure itself.
However, this is not the case for the PS measure. Here, the aggregation we applied is inspired by the window-based pooling and logistic mapping used inside PESQ Rix et al. (2001). Again, considering only time frame indices in $\mathcal{F}^{l}_{i}$ and dropping the rest, let us consider a window of size $W$ frames that slides across the PS measure with a hop size of $H$ frames. Using the $p$-norm, we define the following:
We then calculate the following root mean square expression:
and eventually the aggregated PS measure is given by:
where the constants were chosen according to ITU-T (2007). Here, we penalize lower scores explicitly using the $p$-norm to better match human perceptual aggregation.
The parameters we used for different scenarios and measures are summarized in Table 1.
[…]
PS English 1 1 PESQ
PM English 1 1 Average
PS Spanish 1 1 PESQ
PM Spanish 1 1 Average
PS Music (no drums) 1 1 PESQ
PM Music (no drums) 1 1 Average
PS Music (drums) 0 1 PESQ
PM Music (drums) 0 1 Average
Table 1: The parameters we pre-defined across scenarios for the PS and PM measures.
4.3 Metrics
The preliminary approach we use to evaluate the PM and PS measures is the correlation between their aggregated, utterance-level values and the MOS of the human listeners given in the SEBASS database, which are also utterance-based. We report on two types of correlation coefficients.
$\mathbf{v}^{l}_{i}=$ $\left(v^{1,l}_{i},\ldots,v^{Q,l}_{i}\right)^{T}.$ (44)
Pearson product-moment correlation coefficient (PCC) assumes a linear relationship and therefore captures both rank and spacing between paired observations. High PCC indicates that the metric not only orders systems correctly but also preserves the perceptual distance listeners perceive Benesty et al. (2009). The PCC is measured twice, for the PS and the PM, as follows:
The Spearman rank-order correlation coefficient (SRCC) measures the strength of any monotonic relationship between two variables. Because it depends only on rank, SRCC is insensitive to strictly increasing, possibly non-linear re-scalings of either variable, making it ideal for answering the question “do the PS and PM measures rank systems exactly as listeners do ?” Sedgwick (2014). Let $\mathcal{R}:\mathbb{R}^{Q}\rightarrow R^{Q}$ be the ranking operator, which in the presence of ties assigns the average ranks. Again, the SRCC is measured for the PS and the PM:
We report these correlation coefficients per English, Spanish, and music mixtures scenarios separately. Given a scenario, we assume that it contains $\mathcal{L}$ independent trials such that $l\in\{1,\ldots,\mathcal{L}\}$} with $N^{l}_{\textrm{max}}$ speakers in each, where:
$N^{l}_{\textrm{max}}=\max_{f\in\mathcal{F}^{l}}{N^{l}_{f}}.$ (49)
Architecture Checkpoint Scenario Trans. Layers
WavLM Large microsoft/wavlm-large English 24
WavLM Base microsoft/wavlm-base English 12
Wav2Vec 2.0 Large facebook/wav2vec2-large-lv60 English 24
Wav2Vec 2.0 Base facebook/wav2vec2-base English 12
HuBERT Large facebook/hubert-large-ll60k English 24
HuBERT Base facebook/hubert-base-ls960 English 12
Wav2Vec 2.0 Large facebook/wav2vec2-large-xlsr-53 Spanish 24
AST MIT/ast-finetuned-audioset-10-10-0.4593 Music 12

### v2

@@ 4.3 The Perceptual Match (PM) Measure
where $\hat{k}^{(d)}_{i},\hat{\theta}^{(d)}_{i}$ are well-defined by design for $N_{p}\geq 1$ and a higher score is better. If the output $\hat{a}_{i}^{(d)}$ lies well within the bulk of its distortion cluster, the Gamma-tail probability is near 1, which may indicate a strong perceptual match. As $\hat{a}_{i}^{(d)}$ drifts away, the score decays smoothly toward zero, reflecting degradation. When distortions are tightly concentrated and $\hat{k}^{(d)}_{i}$ or $\hat{\theta}^{(d)}_{i}$ lower, even small mismatches in $\hat{a}_{i}^{(d)}$ lower PM sharply. As $\hat{k}^{(d)}_{i}$ and $\hat{\theta}^{(d)}_{i}$ grow, the PM tolerates larger $\hat{a}_{i}^{(d)}$ deviations.
5 Error Guarantees for the PS and PM Measures
Standing notation and assumptions.
We fix frame $f$ with generally $N_{f}\geq 2$ (1). For this proof, consider the specific case of $N_{f}=2$ with indices $i,j\in S_{f}$. Consider source index $j$ and set m=N1dm=N-1-d as the dimension of the omitted space in the diffusion maps process, so the embedding notations in the retained, omitted, and complete N1{N-1}-dimensional spaces are respectively $\bm{\Psi}^{(d)}_{t}(\mathbf{x}_{j}),\bm{\Psi}^{(m)}_{t}(\mathbf{x}_{j}),\bm{\Psi}_{t}(\mathbf{x}_{j})$ (10). Similarly, clusters $\mathcal{C}_{j}^{(d)}$, $\mathcal{C}_{j}^{(m)}$, and $\mathcal{C}_{j}$ are formed as in $\S$4.1, with means and covariances pairs $\left(\bm{\mu}_{j}^{(d)},\bm{\Sigma}_{j}^{(d)}\right)$, $\left(\bm{\mu}_{j}^{(m)},\bm{\Sigma}_{j}^{(m)}\right)$, and $\left(\bm{\mu}_{j},\bm{\Sigma}_{j}\right)$, and cross-covariance $\bm{C}_{j}\in\mathbb{R}^{d\times m}$. We have empirically prevented ill-conditioning, as all matrix inversions are Tikhonov-regularized (Tikhonov and Arsenin, 1977) with $\epsilon I$, where $\epsilon=10^{-6}$ and $I$ the identity matrix, with context-dependent dimension. When we quantify sampling uncertainty, we use dependence-adjusted effective sample sizes via Bartlett method (Bartlett, 1946), and sub-Gaussian tails for quadratic forms via the dependent Hanson–Wright inequalities (Adamczak, 2015; Vershynin, 2024).
Schur decomposition of full versus truncated Mahalanobis distances. For radius error calculations, from (24) to (36), we assume access to clusters statistics. For the output embedding of source $i$ against cluster $j$, define $\bm{\Delta}^{(d)}_{i,j}=\bm{\Psi}^{(d)}_{t}(\hat{\mathbf{x}}_{i})-\bm{\mu}^{(d)}_{j}$ and $\bm{\Delta}^{(m)}_{i,j}=\bm{\Psi}^{(m)}_{t}(\hat{\mathbf{x}}_{i})-\bm{\mu}^{(m)}_{j}$. The full cluster statistics aggregate as:
$\bm{\mu}_{j}=\begin{bmatrix}\bm{\mu}_{j}^{(d)}\\ \bm{\mu}_{j}^{(m)}\end{bmatrix},\quad\bm{\Delta}_{i,j}=\begin{bmatrix}\bm{\Delta}_{i,j}^{(d)}\\ \bm{\Delta}_{i,j}^{(m)}\end{bmatrix},\quad\bm{\Sigma}_{j}=\begin{bmatrix}\bm{\Sigma}_{j}^{(d)}&\bm{C}_{j}\\ \bm{C}_{j}^{T}&\bm{\Sigma}_{j}^{(m)}\end{bmatrix}$. (24)
Block inversion to (16) via the Schur complement (Horn and Johnson, 2013) yields:
$d_{M}^{2}\!\left(\bm{\Psi}_{t}(\hat{\mathbf{x}}_{i});\bm{\mu}_{j},\bm{\Sigma}_{j}\right)=\underbrace{\left(\bm{\Delta}_{i,j}^{(d)}\right)^{T}\!\left(\bm{\Sigma}_{j}^{(d)}+\epsilon I^{(d)}\right)^{-1}\!\bm{\Delta}_{i,j}^{(d)}}_{:=\,a}+\underbrace{\bm{r}_{i,j}^{T}\bm{S}_{j}^{-1}\bm{r}_{i,j}}_{:=\,b},$ (25)
$\bm{r}_{i,j}=\bm{\Delta}^{(m)}_{i,j}-\bm{C}_{j}^{T}\!\left(\bm{\Sigma}_{j}^{(d)}+\epsilon I^{(d)}\right)^{-1}\!\bm{\Delta}^{(d)}_{i,j},\quad\bm{S}_{j}=\bm{\Sigma}_{j}^{(m)}-\bm{C}_{j}^{T}\!\left(\bm{\Sigma}_{j}^{(d)}+\epsilon I^{(d)}\right)^{-1}\!\bm{C}_{j}.$ (26)
Since $\forall a,b\geq 0:\big|\sqrt{a+b}-\sqrt{a}\big|\leq\sqrt{b}$ (Rudin, 1976, Ch. 5), we bound truncation error to (25):
$|\delta_{i,j}|:=\left|d_{M}\!\left(\bm{\Psi}_{t}(\hat{\mathbf{x}}_{i});\bm{\mu}_{j},\bm{\Sigma}_{j}\right)-d_{M}^{2}\!\left(\bm{\Psi}^{(d)}_{t}(\hat{\mathbf{x}}_{i});\bm{\mu}^{(d)}_{j},\bm{\Sigma}^{(d)}_{j}\right)\right|\;\leq\;\sqrt{\bm{r}_{i,j}^{T}\bm{S}_{j}^{-1}\bm{r}_{i,j}}.$ (27)
PS radius. Let $A_{i},B_{i}$ be the full-space versions of $A_{i}^{(d)},B_{i}^{(d)}$ in (17). Set $|\delta_{i,i}|:=|A_{i}-A_{i}^{(d)}|$ and $|\delta_{i,j^{\ast}}|:=|B_{i}-B_{i}^{(d)}|$, with $j^{\ast}$ as in (17). We empirically confirmed that truncation introduces only mild changes, i.e., $|\delta_{i,i}|,|\delta_{i,j^{\ast}}|\ll A^{(d)}_{i}+B^{(d)}_{i}$. Thus, a first-order Taylor expansion of $\mathrm{PS}_{i}$, the full-space version of $\mathrm{PS}^{(d)}_{i}$, around $\left(A^{(d)}_{i},B^{(d)}_{i}\right)$, is valid. Ultimately, we drop quadratic components that were found negligible, and use $|\delta_{i,i}|$ and $|\delta_{i,j^{\ast}}|$ inside the Taylor expansion, to yield:
$\big|\mathrm{PS}_{i}-\mathrm{PS}_{i}^{(d)}\big|\;\leq\;\frac{B_{i}^{(d)}\,|\delta_{i,i}|+A_{i}^{(d)}\,|\delta_{i,j^{\ast}}|}{\big(A_{i}^{(d)}+B_{i}^{(d)}\big)^{2}}.$ (28)
Combining with (27), the deterministic PS error radius is:
$\big|\mathrm{PS}_{i}-\mathrm{PS}_{i}^{(d)}\big|\leq\frac{B_{i}^{(d)}\sqrt{\bm{r}_{i,i}^{T}\bm{S}_{i}^{-1}\bm{r}_{i,i}}+A_{i}^{(d)}\sqrt{\bm{r}_{i,j^{\ast}}^{T}\bm{S}_{j^{\ast}}^{-1}\bm{r}_{i,j^{\ast}}}}{\big(A_{i}^{(d)}+B_{i}^{(d)}\big)^{2}}.$ (29)
We notice that large residual spread $\bm{\Sigma}_{j}^{(m)}$ or cross-block coupling $\bm{C}_{j}$ inflate (29) through $\bm{S}_{j}^{-1}$.
PM radius.
For source $i$ and every distortion index $p\in\{1,\ldots,N_{p}\}$, we center cluster coordinates at the reference $\mathbf{x}_{i}$, so $\bm{\Delta}^{(d)}_{i,p}=\bm{\Psi}^{(d)}_{t}(\mathbf{x}_{i,p})-\bm{\Psi}^{(d)}_{t}(\mathbf{x}_{i})$ and $\bm{\Delta}^{(m)}_{i,p}=\bm{\Psi}^{(m)}_{t}(\mathbf{x}_{i,p})-\bm{\Psi}^{(m)}_{t}(\mathbf{x}_{i})$. By repeating (24) for the reference-free clusters in the $d$, $m$, and N1N-1 spaces, we obtain:
$\widetilde{\bm{\Sigma}}_{i}=\begin{bmatrix}\widetilde{\bm{\Sigma}}_{i}^{(d)}&\widetilde{\bm{C}}_{i}\\ \widetilde{\bm{C}}_{i}^{T}&\widetilde{\bm{\Sigma}}_{i}^{(m)}\end{bmatrix}$, (30)
where $\widetilde{\bm{\Sigma}}_{i}^{(d)}$ is defined in (19). Exactly as in (25)-(26), we use Schur complement:
$d_{M}^{2}\left(\bm{\Psi}_{t}(\mathbf{x}_{i,p});\bm{\Psi}_{t}(\mathbf{x}_{i}),\widetilde{\bm{\Sigma}}_{i}\right)=\left(\bm{\Delta}_{i,p}^{(d)}\right)^{T}\left(\widetilde{\bm{\Sigma}}_{i}^{(d)}+\epsilon I)^{-1}\right)\bm{\Delta}_{i,p}^{(d)}+\bm{r}_{i,p}^{T}\bm{S}_{i}^{-1}\bm{r}_{i,p},$ (31)
$\bm{r}_{i,p}=\bm{\Delta}^{(m)}_{i,p}-\widetilde{\bm{C}}_{i}^{T}\left(\widetilde{\bm{\Sigma}}_{i}^{(d)}+\epsilon I\right)^{-1}\bm{\Delta}^{(d)}_{i,p},\quad\bm{S}_{i}=\widetilde{\bm{\Sigma}}_{i}^{(m)}-\widetilde{\bm{C}}_{i}^{T}\left(\widetilde{\bm{\Sigma}}_{i}^{(d)}+\epsilon I\right)^{-1}\widetilde{\bm{C}}_{i}.$ (32)
Let $\mathcal{G}_{i}$ be the set of the squared distances in (31) over $p$ and $\mathcal{G}_{i}^{(d)}$ its $d$-dimensional analogue (20). Define per-sample truncation gaps $\delta_{\mathcal{G}_{i},p}:=\bm{r}_{i,p}^{T}\bm{S}_{i}^{-1}\bm{r}_{i,p}\geq 0$ and $\delta_{\max}=\max_{p}\delta_{\mathcal{G}_{i},p}$. Employing elementary algebra and Cauchy–Schwarz inequality (Vershynin, 2024), we obtain the relations (21):
$\left|\mu_{\mathcal{G}_{i}}-\mu_{\mathcal{G}_{i}^{(d)}}\right|=\frac{1}{N_{p}}\sum_{p=1}^{N_{p}}\delta_{\mathcal{G}_{i},p},\quad\left|\sigma^{2}_{\mathcal{G}_{i}}-\sigma^{2}_{\mathcal{G}_{i}^{(d)}}\right|\leq\frac{N_{p}}{N_{p}-1}\!\left(2\delta_{\max}\left(\sigma_{\mathcal{G}_{i}}+\sigma_{\mathcal{G}_{i}^{(d)}}\right)+\delta_{\max}^{2}\right).$ (33)
Again, simple algebra bounds the Gamma-matching parameters (22), with constants $C_{1},C_{2}>0$:
$\left|k_{i}-k_{i}^{(d)}\right|\leq C_{1}\,\delta_{\max}\frac{N_{p}}{N_{p}-1}\frac{\mu_{\mathcal{G}_{i}}+\mu_{\mathcal{G}_{i}^{(d)}}}{\sigma_{\mathcal{G}_{i}^{(d)}}^{2}},\quad\left|\theta_{i}-\theta_{i}^{(d)}\right|\leq C_{2}\,\delta_{\max}\frac{N_{p}}{N_{p}-1}\frac{\sigma_{\mathcal{G}_{i}}^{2}+\sigma_{\mathcal{G}_{i}^{(d)}}^{2}}{\mu_{\mathcal{G}_{i}^{(d)}}^{2}}.$ (34)
Recalling the distance of the output from its cluster, denoted $a_{i}^{(d)}$ (23), we can define using (32):
$d_{M}^{2}\left(\bm{\Psi}_{t}(\hat{\mathbf{x}}_{i});\bm{\Psi}_{t}(\mathbf{x}_{i}),\bm{\tilde{\Sigma}}_{i}\right)-d_{M}^{2}\left(\bm{\Psi}^{(d)}_{t}(\hat{\mathbf{x}}_{i});\bm{\Psi}^{(d)}_{t}(\mathbf{x}_{i}),\bm{\tilde{\Sigma}}^{(d)}_{i}\right)=\bm{r}_{i,a}^{T}\bm{S}_{i}^{-1}\bm{r}_{i,a}:=\delta_{\mathcal{G}_{i},a}.$ (35)
In the full space, $\mathrm{PM}_{i}=Q(k_{i},a_{i}/\theta_{i})$ (23). Its derivatives with respect to its variables are standard and bounded on compact sets (NIST, 2024). Let the truncation ellipsoid $\mathcal{B}_{i}$ be such set, so that $\mathcal{B}_{i}=\Big\{(k^{\prime}_{i},\theta^{\prime}_{i},a^{\prime}_{i}):\Big|k^{\prime}_{i}-k_{i}^{(d)}\Big|\leq\delta_{\mathcal{G}_{i},k},\Big|\theta^{\prime}_{i}-\theta_{i}^{(d)}\Big|\leq\delta_{\mathcal{G}_{i},\theta},\Big|a^{\prime}_{i}-a_{i}^{(d)}\Big|\leq\delta_{\mathcal{G}_{i},a}\Big\}$, with $\delta_{\mathcal{G}_{i},k},\delta_{\mathcal{G}_{i},\theta}$ denoting the bounds in (34). Since $Q$ increases in $k$ and decreases in $x=a/\theta$ for $k,x>0$ (NIST, 2024), the maximum deviation over $\mathcal{B}_{i}$ occurs at a corner, and the radius can be obtained by:
$\left|\mathrm{PM}_{i}-\mathrm{PM}_{i}^{(d)}\right|\leq\max_{(k_{c},\theta_{c},a_{c})\in\partial\mathcal{B}_{i}}\left|\,Q(k_{c},a_{c}/\theta_{c})-Q(k_{i}^{(d)},a_{i}^{(d)}/\theta_{i}^{(d)})\,\right|.$ (36)
Dependence-adjusted sample size. For any cluster $\mathcal{C}_{j}^{(d)}$ with $n_{j}$ dependent points, we use $n_{j,\mathrm{eff}}:=n_{j}\big(1+2\sum_{\ell=1}^{L_{j}}\hat{\rho}_{j,\ell}\big)^{-1}$} with $L_{j}=\min\{\ell:\ |\hat{\rho}_{j,\ell}|<z_{0.975}/\sqrt{n_{j}-\ell}\}$ (Bartlett, 1946).
PS tail bound. We now resort to the retained $d$-dimensional space, and estimate cluster statistics due the finite number of $n_{j,\mathrm{eff}}$ samples. Let $\widehat{\bm{\mu}}_{j}^{(d)}$ and $\widehat{\bm{\Sigma}}_{j}^{(d)}$ be the empirical cluster statistics. Vector and matrix-Bernstein and dependent Hanson–Wright theories (Vershynin, 2024, Props. 2.8.1, 4.7.1), (Adamczak, 2015, Thm. 2.5) give for $\delta_{j,\mu}^{\mathrm{PS}},\delta_{j,\Sigma}^{\mathrm{PS}}\in\left(0,1/2\right)$, with least probabilities $1-\delta_{j,\mu}^{\mathrm{PS}},1-\delta_{j,\Sigma}^{\mathrm{PS}}$:
$\left\|\bm{\mu}_{j}^{(d)}-\widehat{\bm{\mu}}_{j}^{(d)}\right\|_{2}$ $\leq\sqrt{2\,\lambda_{\max}\left(\widehat{\bm{\Sigma}}_{j}^{(d)}\right)\,\ln(2/\delta_{j,\mu}^{\mathrm{PS}})/n_{j,\mathrm{eff}}}=:\Delta_{j,\mu},$ (37)
$\left\|\bm{\Sigma}_{j}^{(d)}-\widehat{\bm{\Sigma}}_{j}^{(d)}\right\|_{2}$ $\leq C\,\lambda_{\max}\left(\widehat{\bm{\Sigma}}_{j}^{(d)}\right)r_{j}+\ln(2/\delta_{j,\Sigma}^{\mathrm{PS}})/n_{j,\mathrm{eff}}=:\Delta_{j,\Sigma},$ (38)
with $r_{j}=\mathrm{trace}(\widehat{\bm{\Sigma}}_{j}^{(d)})/\lambda_{\max}\left(\widehat{\bm{\Sigma}}_{j}^{(d)}\right)$. A first-order perturbation of $A_{i}^{(d)},B_{i}^{(d)}$ (17) yields the bounds:
$\varepsilon^{\mathrm{PS}}\left(\widehat{A}_{i}^{(d)}\right)\leq 2\sqrt{\widehat{A}_{i}^{(d)}}\Delta_{i,\bm{\mu}}\sqrt{\lambda_{\textrm{max}}\left(\widehat{\mathbf{\Sigma}}_{i}^{(d)}\right)/\tilde{\lambda}_{\textrm{min}}\left(\widehat{\mathbf{\Sigma}}_{i}^{(d)}\right)}+\widehat{A}_{i}^{(d)}\Delta_{i,\mathbf{\Sigma}}/\lambda_{\textrm{max}}\left(\widehat{\mathbf{\Sigma}}_{i}^{(d)}\right),$ (39)
$\varepsilon^{\mathrm{PS}}\left(\widehat{B}_{i}^{(d)}\right)\leq 2\sqrt{\widehat{B}_{i}^{(d)}}\Delta_{j^{\ast},\bm{\mu}}\sqrt{\lambda_{\textrm{max}}\left(\widehat{\mathbf{\Sigma}}_{j^{*}}^{(d)}\right)/\tilde{\lambda}_{\textrm{min}}\left(\widehat{\mathbf{\Sigma}}_{j^{*}}^{(d)}\right)}+\widehat{B}_{i}^{(d)}\Delta_{j^{\ast},\mathbf{\Sigma}}/\lambda_{\textrm{max}}\left(\widehat{\mathbf{\Sigma}}_{j^{\ast}}^{(d)}\right).$ (40)
With $L_{i}^{\mathrm{PS}}$ being the Euclidean gradient norm of $\mathrm{PS}_{i}^{(d)}$ at $(A_{i}^{(d)},B_{i}^{(d)})$, for $\delta_{i}^{\mathrm{PS}}=\delta_{i,\mu}^{\mathrm{PS}}+\delta_{i,\Sigma}^{\mathrm{PS}}\in\left(0,1\right)$
$\left|\widehat{\mathrm{PS}}_{i}^{(d)}-\mathrm{PS}_{i}^{(d)}\right|\leq L_{i}^{\mathrm{PS}}\sqrt{\varepsilon^{\mathrm{PS}}\left(\widehat{A}_{i}^{(d)})+\varepsilon^{\mathrm{PS}}(\widehat{B}_{i}^{(d)}\right)}\quad\text{w.p.\ }\geq 1-\delta_{i}^{\mathrm{PS}}.$ (41)
PM tail bound. Let $R_{i}=\max_{g\in\mathcal{G}_{i}}g$ and choose confidence levels $\delta_{i,\mu}^{\mathrm{PM}},\delta_{i,\sigma}^{\mathrm{PM}},\delta_{i,a}^{\mathrm{PM}}\in(0,1/3)$. Concentration bounds for the output quadratic form via Hanson–Wright (Vershynin, 2024, Props. 2.8.1, 4.7.1) yields
$\big|\mu_{\mathcal{G}_{i}^{(d)}}-\widehat{\mu}_{\mathcal{G}_{i}^{(d)}}\big|\leq\sqrt{\frac{2\widehat{\sigma}_{\mathcal{G}_{i}^{(d)}}^{2}\ln(2/\delta_{i,\mu}^{\mathrm{PM}})}{N_{p}}}+\frac{3R_{i}\ln(2/\delta_{i,\mu}^{\mathrm{PM}})}{N_{p}},$ (42)
$\big|\sigma_{\mathcal{G}_{i}^{(d)}}-\widehat{\sigma}_{\mathcal{G}_{i}^{(d)}}\big|\leq\sqrt{\frac{2R_{i}^{2}\ln(2/\delta_{i,\sigma}^{\mathrm{PM}})}{N_{p}}}+\frac{3R_{i}^{2}\ln(2/\delta_{i,\sigma}^{\mathrm{PM}})}{N_{p}},\,\big|a_{i}^{(d)}-\widehat{a}_{i}^{(d)}\big|\leq R_{i}\sqrt{\frac{\ln(2/\delta_{i,a}^{\mathrm{PM}})}{N_{p}}}.$ (43)
Just like in (39) and (40), these are mapped to bounds on the parameters $\left(k,\theta,a\right)$, yielding $\Delta_{i,k},\Delta_{i,\theta},\Delta_{i,a}$. Consider $\delta_{i}^{\mathrm{PM}}=\delta_{i,\mu}^{\mathrm{PM}}+\delta_{i,\sigma}^{\mathrm{PM}}+\delta_{i,a}^{\mathrm{PM}}\in(0,1)$, then for the local box $\mathcal{B}_{i}^{\mathrm{loc}}=\{\,|k-k_{i}^{(d)}|\leq\Delta_{i,k},\;|\theta-\theta_{i}^{(d)}|\leq\Delta_{i,\theta},\;|a-a_{i}^{(d)}|\leq\Delta_{i,a}\,\}$}:
$\big|\widehat{\mathrm{PM}}_{i}^{(d)}-\mathrm{PM}_{i}^{(d)}\big|\leq\max_{(k_{c},\theta_{c},a_{c})\in\partial\mathcal{B}_{i}^{\mathrm{loc}}}\big|\,Q(k_{c},a_{c}/\theta_{c})-Q(\widehat{k}_{i}^{(d)},\widehat{a}_{i}^{(d)}/\widehat{\theta}_{i}^{(d)})\,\big|\quad\text{w.p.\ }\geq 1-\delta_{i}^{\mathrm{PM}}.$ (44)
@@ 6.1 Database
We use the Subjective Evaluation of Blind Audio Source Separation (SEBASS) database (Kastner and Herre, 2022), a public collection of expertly curated listening tests that aggregates $11,000$ ratings for more than $900$ separated signals across five evaluation campaigns. SEBASS covers speech mixtures of 4 male or 4 female speakers, each consisting of English and Spanish pairs. As realistic conversations are monolingual, we separate each mixture into English and Spanish speakers pairs. Also included are music mixtures with drums and without drums, each with 3 sources. Namely, $N_{f}\in\{2,3\}$ (1). The split between drum and no-drum mixtures is crucial, as percussion transients create perceptual and algorithmic masking distinct from harmonic content. Each mixture was processed by 32 source separation systems, ranging from classic approaches to deep-learning models. Outputs with 10 s duration, sampled at 16 kHz, were judged by 15 certified raters under the MUSHRA standard (Schoeffler et al., 2018), which grades output quality between 0 and $100$ relative to a reference.
6.2 Pre-processing and Performance Evaluation
SEBASS provides MOS values at the utterance level. Since our PM and PS measures operate at much finer temporal resolutions, with frame sizes of $L=400$ for speech and $L=324$ samples for music (1), aggregation from the frame-level to the utterance-level is required to enable comparison with human MOS. PM values are aggregated using a simple average, while PS values are aggregated with a perceptually-weighted scheme inspired by PESQ. For performance evaluation, we correlate the aggregated PM and PS values with the utterance-level MOS values using the Pearson product-moment correlation coefficient (PCC) (Benesty et al., 2009) and the Spearman rank-order correlation coefficient (SRCC) (Sedgwick, 2014). We set $\alpha=1$ (3) to eliminate density-dependent bias from the embedding, and $t=1$ (6) to keep the diffusion operator focused on local structures. The retained dimension $d$ is in $\left[20,40\right]$ (10), using $\tau=0.99$ (11), as done on (FjellstromNyström2022).